# Optimal Stopping in Latent Diffusion Models

### v1

@@ 3 Notations and problem setup
$P_{d}\overleftarrow{\hat{X}}_{t}\sim\mathcal{N}(0,a_{t}^{2}I_{d}+b_{t}^{2}P_{d}\hat{\Sigma}P_{d}^{\top}).$
In practical applications, some numerical scheme is used to solve the backward SDE (5). Typical choices include replacing the initial distribution $(P_{d})_{\#}\hat{p}_{T}$ in the SDE (5) by a standard Gaussian distribution $\mathcal{N}(0,I_{D})$, which is a valid approximation when $T$ is large. In addition, for numerical stability, the backward diffusion is early stopped at a preset time $T-\delta$ (see, e.g., Yang et al., 2023).
We quantify the distance between distributions by the Wasserstein-2 distance (Villani, 2008), which is equivalent in the Gaussian case to the Fréchet distance (Heusel et al., 2017):
$d_{F}^{2}(\mathcal{N}(\mu_{1},\Sigma_{1}),\mathcal{N}(\mu_{2},\Sigma_{2}))=\|\mu_{1}-\mu_{2}\|_{2}^{2}+\textnormal{tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{2}^{1/2}\Sigma_{1}\Sigma_{2}^{1/2})^{1/2}).$ (6)
With a small abuse of notation, we use $d_{F}(X,Y)$ to denote the distance between the distributions of the random variables $X$ and $Y$.
4 Optimal dimension reduction and stopping time
In this section, we address the important question of how dimensionality reduction affects the diffusion process with respect to the intrinsic geometric structure of the data. Our analysis focuses on selecting the rank $d$ of the projection matrix $P_{d}$ and the stopping time of the diffusions (2) and (5).
4.1 An analysis of non-monotonic behavior of Fréchet distance
Proposition 1.
Let $P_{d}\overleftarrow{X}_{t}$ and $P_{d}\overleftarrow{\hat{X}}_{t}$ be given as in (2) and (5), respectively. For $d\in\{1,\ldots,D\}$, the Fréchet distance $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X}_{0})$ is non-increasing with respect to $t$. On the other hand, $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X}_{0})$ is non-increasing if and only if
Roughly speaking, the variance of each backward diffusion component $P_{d}\overleftarrow{\hat{X}}_{t}$ scales monotonically from an initial value close to 1 ($a_{T}^{2}+b_{T}^{2}\hat{\sigma}_{d^{\prime}}^{2}\approx 1$) to the estimated variance $\hat{\sigma}_{d^{\prime}}^{2}$ or any values satisfying (7), e.g., when $\hat{\sigma}_{d^{\prime}}^{2}$ are given by an oracle. The distance $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ is therefore minimized when the process variance is closest to the true set of variances $(\sigma_{d^{\prime}}^{2})_{1\leq d^{\prime}\leq d}$, which happens before time $T$ under condition (7).
$\sum_{d^{\prime}=1}^{D}(1-\frac{\sigma_{d^{\prime}}}{\hat{\sigma}_{d^{\prime}}})(1-\hat{\sigma}_{d^{\prime}}^{2})=\sum_{d^{\prime}=1}^{d_{0}}(1-\frac{\sigma_{d^{\prime}}}{\hat{\sigma}_{d^{\prime}}})(1-\hat{\sigma}_{d^{\prime}}^{2})+D-d_{0}.$
For large enough $n$ and with high probability, $|\sigma_{d^{\prime}}-\hat{\sigma}_{d^{\prime}}|\leq\hat{\sigma}_{d^{\prime}}$ for every $d^{\prime}\in\{1,\ldots,d_{0}\}$, thus
$\sum_{d^{\prime}=1}^{D}(1-\frac{\sigma_{d^{\prime}}}{\hat{\sigma}_{d^{\prime}}})(1-\hat{\sigma}_{d^{\prime}}^{2})\geq\sum_{d^{\prime}=1}^{d_{0}}\big(-1-\max_{d^{\prime}\in\{1,\ldots,d_{0}\}}\hat{\sigma}_{d^{\prime}}^{2}\big)+D-d_{0}=D-\big(2+\max_{d^{\prime}\in\{1,\ldots,d_{0}\}}\hat{\sigma}_{d^{\prime}}^{2}\big)d_{0}.$
However, if projecting the diffusion onto the $d_{0}$-dimensional linear subspace in which the data distribution lies, the $D-d_{0}$ term in the computation above vanishes, and we are left with the sum up to $d_{0}$. Then the behavior of the Fréchet distance is linked to how the model estimates the variances of the data. If, for most $d^{\prime}$, the sign of $1-\sigma_{d^{\prime}}/\hat{\sigma}_{d^{\prime}}$ matches the sign of $1-\hat{\sigma}_{d^{\prime}}^{2}$, the Fréchet distance exhibits monotonic behavior. Conversely, if most of the signs differ, the Fréchet distance is non-monotonic.
Figure 3: $\bar{a}^{-2}$ in the Ornstein-Uhlenbeck process.
4.2 Optimal projection at time $t$
In this subsection, we continue the study of the interaction of the dimension of projection and the stopping time. In contrast to the previous subsection, we show that for each fixed time $t$, there exists an optimal projection $P_{d}$. We still consider Gaussian data with independent components (3). Recall that $a$ is defined in (4) and that it is an increasing map from $[0,T]$ to $0,a_{T}]$. We then let $\bar{a}^{-2}:\mathbb{R}\cup\{\infty\}\to[0,T]$ be the extended inverse function of $a^{2}$ (see plot in Figure 3), meaning that
$t_{d}=T-\bar{a}^{-2}\Big(\frac{3\sigma_{d}^{2}}{(1-\sigma_{d}^{2})_{+}}\Big)\quad\text{and}\quad\hat{t}_{d}=T-\bar{a}^{-2}\Big(\frac{4\sigma_{d}^{2}-\hat{\sigma}_{d}^{2}}{(1-\hat{\sigma}_{d}^{2})_{+}}\Big).$
Proposition 2.
$d_{F}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X_{0}})=\min_{d^{\prime}\in\{1,\ldots,D\}}d_{F}(P_{d^{\prime}}^{\top}P_{d^{\prime}}\overleftarrow{X}_{t},\overrightarrow{X_{0}}).$
$d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})=\min_{d^{\prime}\in\{1,\ldots,D\}}d_{F}(P_{d^{\prime}}^{\top}P_{d^{\prime}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
Figure 4: Plots of $d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ as a function of the diffusion time $t$, for two sets of variances $(\sigma_{1},\dots,\sigma_{D})$. (left) All the $\sigma_{i}$ are nonzero. As expected from Proposition 2, the $d$-dimensional projection is optimal in $t_{d},t_{d+1})$. (right) The data is supported on a linear subspace of dimension $d_{0}=4$ with $D=6$. As expected from Proposition 3, we observe that the minimum distance is achieved in dimension $d_{0}$ and with early stopping. LogSNR in the $x$-axis is a remapping of time $t$, defined as $\log(b_{t}^{2}/a_{t}^{2})$, which we use to increase readability. Experimental details are in Appendix D.
Proposition 2 quantifies a direct link between early stopping in the backward diffusion process and dimensionality reduction. It reveals a time-dependent trade-off: distributions at early stages of the backward process are best approximated in lower-dimensional spaces, while higher dimensions become necessary to faithfully reconstruct the data as $t\to T$, as illustrated in Figure 4 (left). In other words, at an early time step, projecting onto an unnecessarily high-dimensional space can introduce more noise than signal, making a lower-dimensional representation more accurate. Notice that when $4\sigma_{d}^{2}\geq 1$, both $t_{d}$ and $\hat{t}_{d}$ are equal to 0. This implies that a component whose variance is sufficiently large should always be included in the projection, aligning with the intuition that major components are essential for representation. These results hold for backward processes using scores based on either the true or empirical variances.
We next characterize the behavior of the optimal latent dimension and stopping time when data lies on a $d_{0}$-dimensional subspace, providing a similar result to Proposition 2. This analysis allows us to precisely determine these two key parameters, as shown next.
Proposition 3.
Assume that $\Sigma=\textnormal{diag}(\sigma^{2},\ldots,\sigma^{2},0,\ldots,0)$ with the last $D-d_{0}$ entries equal to 0. Let $\varepsilon\in(0,1)$. Then, there exists $\hat{\delta}_{d_{0}}\in[0,T]$ such that with probability $1-2d_{0}e^{-\frac{n}{8}}$,
$d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{T-\hat{\delta}_{d_{0}}},\overrightarrow{X_{0}})=\min_{\begin{subarray}{c}t\in[0,T]\\ d^{\prime}\in\{1,\ldots,D\}\end{subarray}}d_{F}(P_{d^{\prime}}^{\top}P_{d^{\prime}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
The proposition shows that the optimal generation strategy for data with a low-rank structure involves both early stopping and projection (see Figure 4 (right) for an illustration). The proof indicates that, under the non-monotonicity condition of Proposition 1, the optimal early stopping time $T-\hat{\delta}_{d_{0}}$ is strictly before $T$. Beyond preventing numerical instability as $t\to T$ (e.g., Yang et al., 2023), Proposition 3 thus offers a new justification for early stopping. In other words, stopping at a specific time $\hat{\delta}_{d_{0}}$ is not merely a practical fix, but an optimal strategy to improve generation quality by minimizing the distance between the generated and true data distributions.
Furthermore, this result confirms the intuition that confining the generative process to the dataset’s intrinsic dimensionality is the most effective approach for low-rank data. This strategy is not only computationally more efficient than running the diffusion in the ambient space, but also enhances generation quality by avoiding the noise introduced by superfluous dimensions. This suggests a principled guideline for practitioners: identifying and restricting the diffusion process to the data’s intrinsic dimensionality leads to more accurate and robust generative models.
5 Performance of the score matching ERM
$\mathcal{R}(s)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{t\sim\mathcal{T},\varepsilon\sim\mathcal{N}(0,I_{D})}\left\|s(b_{t}X_{i}+a_{t}\varepsilon,t)+\frac{\varepsilon}{a_{t}}\right\|^{2},$ (9)
$\nabla\log p_{t}(x)=-(a_{t}^{2}I_{D}+b_{t}^{2}\Sigma)^{-1}x,$ (10)
$\mathcal{F}_{C}=\left\{s_{M}:\mathbb{R}^{d}\times\mathbb{R}_{+}\to\mathbb{R}^{d}:\right$. $s_{M}(x,t)=-M(t)x,\qquad M(t)=\textnormal{diag}(m_{1}(t),\ldots,m_{D}(t)),$
$\qquad\qquad\qquad\qquad\qquad\quad\left.m_{i}\in\mathcal{L}_{2}(\mathbb{R}_{+},\mathbb{R}),\|m_{i}\|_{\infty}<C\right\}.$
Proposition 4.
Our goal in the following is to characterize the optimal latent dimension when using the score defined by Proposition 4. For this purpose, as before, we quantify the distance between the data distribution and the distribution generated by the backward process for $d\in\{1,\ldots,D\}$. We do not consider early stopping here, to focus on the influence of the regularization parameter $C$ on the choice of the latent dimensionality. For simplicity, we keep the data distribution $p_{0}$ to be a Gaussian distribution with independent component, and specialize to the Ornstein-Uhlenbeck process. In this case, the sample $\overleftarrow{\tilde{X}}_{t}$ are generated by the backward SDE for $t\in[0,T]$
$d\overleftarrow{\tilde{X}}_{t}=(\overleftarrow{\tilde{X}}_{t}+2s_{\hat{M}}(\overleftarrow{\tilde{X}}_{t},T-t))dt+\sqrt{2}d\overleftarrow{W_{t}},\quad\overleftarrow{\tilde{X}}_{0}\sim\mathcal{N}(0,I_{D}).$
Proposition 5.
Define $1\leq d_{1}\leq d_{2}\leq D$ as follows:
$d_{1}=\max\{d^{\prime}\in\{1,\ldots,D\}:1/C\leq\hat{\sigma}_{d^{\prime}}^{2}\}\;\textnormal{and}\;d_{2}=\min\Big\{d^{\prime}\in\{1,\ldots,D\}:\frac{1}{2C-1}\geq 4\sigma_{d^{\prime}}^{2}\Big\}.$
$d_{F}(P_{d_{\min}}^{\top}P_{d_{\min}}\overleftarrow{\tilde{X}}_{T},\overrightarrow{X_{0}})=\min_{d^{\prime}\in\{1,\ldots,D\}}\big\{d_{F}(P_{d^{\prime}}^{\top}P_{d^{\prime}}\overleftarrow{\tilde{X}}_{T},\overrightarrow{X_{0}})\big\}.$
To gain intuition into Proposition 5, let us consider some illustrative cases depending on the weight constraint $C$ (for the full derivation of these cases, see Appendix A.6). First, when $C=\infty$ and the data covariance is non-singular, we get $d_{1}=d_{2}=D$, thus Proposition 5 suggests to take the projection matrix $I_{D}$, which is expected since $\mathcal{F}_{C}$ is then large enough to contain the true empirical score function. Second, consider the scenario when the data distribution lies on a linear subspace of dimension $d_{0}$. If $C$ is large enough, we obtain $d_{1}=d_{2}=d_{0}$, meaning that the projection onto the data subspace is the optimal sampling strategy, which is in line with Proposition 3. Finally, the optimal projection can also be made explicit for exponentially-decaying covariance spectrum.
Corollary 1.
$d_{\min}\in\{d,d+1\}.$
Interestingly, when the covariance structure decays exponentially, the capacity of the score-function class—captured here by the parameter CC—is directly tied to the optimal projection dimension. The latter scales logarithmically in the parameter $C$ since $\hat{\sigma}_{d_{\min}}^{2}=\lambda^{-d_{\min}}\approx 1/C$. An analogous result can be derived for covariance structures with power-law decay.
6 Generalization to arbitrary Gaussian distributions
We now explain how to generalize some of our preceding analysis from Gaussian distributions with diagonal covariance matrices to the more general case $p_{0}=\mathcal{N}(0,\Sigma)$ for arbitrary $\Sigma$, and the backward processes $\overleftarrow{X}_{t}$ and $\overleftarrow{\hat{X}}_{t}$ given as in (2) and (5) with the new general data distribution $p_{0}$. Our goal is to establish a result analogous to Proposition 2, that is, to characterize the optimal latent dimension given a stopping time of the diffusion process.
Proposition 6.
For $2\leq d\leq D$ and $t\in[t_{d},t_{d+1})$, we have
$d_{F}(OP_{d}^{\top}P_{d}O^{\top}\overleftarrow{X}_{t},\overrightarrow{X_{0}})=\min_{d^{\prime}\in\{1,\ldots,D\}}d_{F}(OP_{d^{\prime}}^{\top}P_{d^{\prime}}O^{\top}\overleftarrow{X}_{t},\overrightarrow{X_{0}}).$
However, in practical applications, one rarely has access to the true underlying covariance matrix $\Sigma$ or its eigenbasis $O$. Instead, one must rely on estimations derived from observed data, where PCA is commonly used. Denote $\hat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}X_{i}X_{i}^{\top}$ to be the empirical covariance matrix. Applying a spectral decomposition yields $\hat{\Sigma}=\hat{O}\textnormal{diag}(\hat{\sigma}_{1}^{2},\ldots,\hat{\sigma}_{D}^{2})\hat{O}^{\top}$, where $\hat{O}$ contains the orthonormal eigenvectors and $\hat{\sigma}_{D}^{2}<\ldots<\hat{\sigma}_{1}^{2}$ are the corresponding eigenvalues. Denote $S(\Sigma)=\sum_{d^{\prime}=1}^{D}\max(\sigma_{d},\sigma_{d}^{2})$. For $u\geq 0$ and $d\in\{2,\ldots,D\}$, we let $\hat{T}_{d}(u)$ and $\hat{t}_{d}(u)$ be
$\hat{T}_{d}(u)$ $=T-\bar{a}^{-2}\Bigg(\frac{\hat{\sigma}_{d}^{2}-4S(\Sigma)\varepsilon_{u}+2\hat{\sigma}_{d}\sqrt{\hat{\sigma}_{d}^{2}-4S(\Sigma)\varepsilon_{u}}}{(1-\hat{\sigma}_{d}^{2})_{+}}\Bigg),$
$\hat{t}_{d}(u)$ $=T-\bar{a}^{-2}\Bigg(\frac{\hat{\sigma}_{d}^{2}+4S(\Sigma)\varepsilon_{u}+2\hat{\sigma}_{d}\sqrt{\hat{\sigma}_{d}^{2}+4S(\Sigma)\varepsilon_{u}}}{(1-\hat{\sigma}_{d}^{2})_{+}}\Bigg),$
$0=\hat{T}_{1}(u)<\hat{t}_{2}(u)<\hat{T}_{2}(u)<\cdots<\hat{t}_{D}(u)<\hat{T}_{D}(u)<\hat{t}_{D+1}(u)=T.$
Proposition 7.
$d_{F}(\hat{O}P_{d}^{\top}P_{d}\hat{O}^{\top}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})=\min_{d^{\prime}\in\{1,\ldots,D\}}d_{F}(\hat{O}P_{d^{\prime}}^{\top}P_{d^{\prime}}\hat{O}^{\top}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
This proposition generalizes the result of Proposition 2 to the case of a general Gaussian data distribution. The analysis reveals that, for any latent dimension $d$, there exists a time interval where a $d$-dimensional projected diffusion process minimizes the distance to the target distribution with high probability. Notably, this result is consistent with our previous conclusions. In the idealized scenario where the variance estimation error is zero (i.e., $\varepsilon_{u}=0$, implying $\hat{\Sigma}=\Sigma$,) the formula for the optimal time $\hat{t}_{d}=\hat{T}_{d}$ simplifies precisely to the one derived in Proposition 6.
7 Conclusion
Acknowledgments
P.M. is supported by a Google PhD Fellowship.
Appendix
Appendix A Proofs of results
A.1 Proof of Proposition 1
We show the equivalent statement: $t\in[0,T]\mapsto d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})$ is non-decreasing if and only if (7) holds. We start by calculating the Fréchet distance $d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})$ by using (6):
$d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X}_{0})$ $=\sum_{d^{\prime}=d+1}^{D}\sigma_{d^{\prime}}^{2}+\sum_{d^{\prime}=1}^{d}\Big(b_{t}^{2}\hat{\sigma}_{d^{\prime}}^{2}+a_{t}^{2}+\sigma_{d^{\prime}}^{2}-2\sigma_{d^{\prime}}\sqrt{a_{t}^{2}+b_{t}^{2}\hat{\sigma}_{d^{\prime}}^{2}}\Big)$
@@ A.1 Proof of Proposition 1
Since $t\mapsto a_{t}^{2}$ is strictly increasing with $a_{0}=0$, the monotonicity of $d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})$ with respect to $t$ is equivalent to the monotonicity with respect to $a_{t}^{2}$. By considering the function $f:[0,a_{T}^{2}]\to\mathbb{R}$ defined by
@@
we see that $d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})$ is non-decreasing if and only if $f$ is non-decreasing. Additionally,
@@
$d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})=f(a_{t}^{2})+\sum_{d^{\prime}=d+1}^{D}\sigma_{d^{\prime}}^{2}$
is non-decreasing if and only if $f^{\prime}(0)\geq 0$, i.e., if and only if $\sum_{d^{\prime}=1}^{d}(1-\frac{\sigma_{d^{\prime}}}{\hat{\sigma}_{d^{\prime}}})(1-\hat{\sigma}_{d^{\prime}}^{2})\geq 0$. This shows the second statement of the proposition. The monotonicity of $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{X}_{T-t},\overrightarrow{X_{0}})$ can be shown by replacing $\hat{\sigma}_{d^{\prime}}$ with $\sigma_{d^{\prime}}$ in the derivative $f^{\prime}$, which is 0 when $a_{t}=0$.
A.2 Proof of Proposition 2
The first part of Proposition 2 concerns the minimization of $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X_{0}})$. Recall that $t_{d}=T-\bar{a}^{-2}\Big(\frac{3\sigma_{d}^{2}}{1-\sigma_{d}^{2}}\Big)$. To prove that $P_{d}\overleftarrow{X}_{t}$ achieves the minimal distance to the target for $t\in[t_{d},t_{d+1})$ (where the time interval is fixed), we will demonstrate how the distance $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X_{0}})$ behaves as a function of the projection dimension $d$. Specifically, we aim to show that, for any $d\in\{2,\ldots,D\}$,
$d_{F}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X_{0}})$ $\leq d_{F}(P_{d-1}^{\top}P_{d-1}\overleftarrow{X}_{t},\overrightarrow{X_{0}})\quad\textnormal{iff }t\geq t_{d}.$ (11)
This inequality in turn implies that for a given $t$ in a fixed interval $t_{d},t_{d+1})$, the minimum distance $d_{F}(P_{d}^{\top}P_{d}\overrightarrow{X_{t}},\overrightarrow{X_{0}})$ is attained by the projected process $P_{d}\overrightarrow{X_{t}}$ in dimension $d$.
To establish them, we first explicitly compute the Fréchet distance $d_{F}(P_{d}^{\top}P_{d}\overrightarrow{X_{t}},\overrightarrow{X_{0}})$. Recall that the Fréchet distance between two zero-mean Gaussian distributions $\mathcal{N}(0,\Sigma_{1})$ and $\mathcal{N}(0,\Sigma_{2})$ is given by $\textnormal{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{1}^{1/2}\Sigma_{2}\Sigma_{1}^{1/2})^{1/2})$, and that the covariance matrix of $P_{d}\overleftarrow{X}_{t}$ is equal to $P_{d}(a_{T-t}^{2}I_{d}+b_{T-t}^{2}\Sigma)P_{d}$. Therefore, it is possible to calculate the Fréchet distance to the target for the projected processes directly, as, for any $d\in\{1,\ldots,D\}$,
$d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X}_{0})=\sum_{j=1}^{D}\sigma_{j}^{2}+\sum_{j=1}^{d}(a_{T-t}^{2}+b_{T-t}^{2}\sigma_{j}^{2})-2\sum_{j=1}^{d}\sigma_{j}\sqrt{a_{T-t}^{2}+b_{T-t}^{2}\sigma_{j}^{2}},$
@@ A.2 Proof of Proposition 2
  • otherwise, $\Delta_{d,t}$ is non-positive if and only if $a_{T-t}^{2}\leq\frac{3\sigma_{d}^{2}}{1-\sigma_{d}^{2}}$ which is equivalent to
@@
Putting things together, we obtain that $d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{X}_{t},\overrightarrow{X}_{0})-d_{F}^{2}(P_{d-1}^{\top}P_{d-1}\overleftarrow{X}_{t},\overrightarrow{X}_{0})$ is non-positive iff $t\geq t_{d}$, which is exactly (11).
@@ A.3 Proof of Proposition 3
$d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})=\min_{d^{\prime}\in\{1,\ldots,D\}}d_{F}(P_{d^{\prime}}^{\top}P_{d^{\prime}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
@@
$d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{T-\hat{\delta}_{d_{0}}},\overrightarrow{X_{0}})=\min_{\begin{subarray}{c}t\in[0,T]\\ d^{\prime}\in\{1,\ldots,D\}\end{subarray}}d_{F}(P_{d^{\prime}}^{\top}P_{d^{\prime}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
Let $\varepsilon\in(0,1)$. We first note that according to Proposition 9, by the union bound, with probability $1-2d_{0}e^{-\frac{\epsilon^{2}n}{4(1+\epsilon)}}\geq 1-2d_{0}e^{-\frac{\epsilon^{2}n}{8}}$ we have $|\sigma^{2}-\hat{\sigma}_{d}^{2}|\leq\varepsilon\sigma^{2}$ for all $d\in\{1,\ldots,d_{0}\}$. We work under this event in the remainder of the proof. In particular, for all $d\in\{1,\ldots,d_{0}\}$, $\sigma_{d}^{2}=\sigma^{2}\geq\hat{\sigma}_{1}^{2}/(1+\varepsilon)$. Thus, by separating cases depending on whether $4\sigma_{d}^{2}\leq 1$, a short calculation gives that

### v2

@@ Appendix A Discussion on diffusion and autoencoders
$P\overleftarrow{\hat{X}}_{t}\sim\mathcal{N}(0,a_{T-t}^{2}I_{d}+b_{T-t}^{2}P\hat{\Sigma}P^{\top}).$ (12)
[…]
Appendix B Proofs of results
B.1 Proof of Proposition 4.1
[…]
@@ Appendix B Proofs of results
[…]
@@
[…]
@@
$\Delta_{d,t}(\hat{\Sigma})=f(a_{t}^{2})+\sum_{d^{\prime}=d+1}^{D}\sigma_{d^{\prime}}^{2}$
is non-decreasing if and only if $f^{\prime}(0)\geq 0$, i.e., if and only if $\sum_{d^{\prime}=1}^{d}(1-\frac{\sigma_{d^{\prime}}}{\hat{\sigma}_{d^{\prime}}})(1-\hat{\sigma}_{d^{\prime}}^{2})\geq 0$. This shows the second statement of the proposition. The monotonicity of $\Delta_{d,t}(\Sigma)$ can be shown by replacing $\hat{\sigma}_{d^{\prime}}$ with $\sigma_{d^{\prime}}$ in the derivative $f^{\prime}$, which is 0 when $a_{t}=0$.
B.2 Proof of Proposition 4.2
The first part of Proposition 4.2 concerns the minimization of $\Delta_{d,t}(\Sigma)$. Recall that $t_{d}=T-\bar{a}^{-2}\Big(\frac{3\sigma_{d}^{2}}{1-\sigma_{d}^{2}}\Big)$. To prove that $\Delta_{d,t}(\Sigma)$ achieves the minimum for $t\in[t_{d},t_{d+1})$ (where the time interval is fixed), we will demonstrate how the distance $\Delta_{d,t}(\Sigma)$ behaves as a function of the projection dimension $d$. Specifically, we aim to show that, for any $d\in\{2,\ldots,D\}$,
$\Delta_{d,t}(\Sigma)$ $\leq\Delta_{d-1,t}(\Sigma)\quad\textnormal{iff }t\geq t_{d}.$ (13)
This inequality in turn implies that for a given $t$ in a fixed interval $t_{d},t_{d+1})$, the minimum distance $\Delta_{d,t}(\Sigma)=d_{F}^{2}(P_{d}^{\top}P_{d}\overrightarrow{X_{t}},\overrightarrow{X_{0}})$ is attained by the projected process $P_{d}\overrightarrow{X_{t}}$ in dimension $d$.
To establish them, we first explicitly compute the Fréchet distance $\Delta_{d,t}(\Sigma)$. Recall that the Fréchet distance between two zero-mean Gaussian distributions $\mathcal{N}(0,\Sigma_{1})$ and $\mathcal{N}(0,\Sigma_{2})$ is given by $\textnormal{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{1}^{1/2}\Sigma_{2}\Sigma_{1}^{1/2})^{1/2})$, and that the covariance matrix of $P_{d}\overleftarrow{X}_{t}$ is equal to $P_{d}(a_{T-t}^{2}I_{d}+b_{T-t}^{2}\Sigma)P_{d}$. Therefore, it is possible to calculate the Fréchet distance to the target for the projected processes directly, as, for any $d\in\{1,\ldots,D\}$,
$\Delta_{d,t}(\Sigma)=\sum_{j=1}^{D}\sigma_{j}^{2}+\sum_{j=1}^{d}(a_{T-t}^{2}+b_{T-t}^{2}\sigma_{j}^{2})-2\sum_{j=1}^{d}\sigma_{j}\sqrt{a_{T-t}^{2}+b_{T-t}^{2}\sigma_{j}^{2}},$
@@
  • otherwise, $\textnormal{diff}_{d,t}$ is non-positive if and only if $a_{T-t}^{2}\leq\frac{3\sigma_{d}^{2}}{1-\sigma_{d}^{2}}$ which is equivalent to
@@
Putting things together, we obtain that $\Delta_{d,t}(\Sigma)-d_{F}^{2}(P_{d-1}^{\top}P_{d-1}\overleftarrow{X}_{t},\overrightarrow{X}_{0})$ is non-positive iff $t\geq t_{d}$, which is exactly (13).
@@
[…]
@@
[…]
Let $\varepsilon\in(0,1)$. We first note that according to Proposition C.1, by the union bound, with probability $1-2d_{0}e^{-\frac{\epsilon^{2}n}{4(1+\epsilon)}}\geq 1-2d_{0}e^{-\frac{\epsilon^{2}n}{8}}$ we have $|\sigma^{2}-\hat{\sigma}_{d}^{2}|\leq\varepsilon\sigma^{2}$ for all $d\in\{1,\ldots,d_{0}\}$. We work under this event in the remainder of the proof. In particular, for all $d\in\{1,\ldots,d_{0}\}$, $\sigma_{d}^{2}=\sigma^{2}\geq\hat{\sigma}_{1}^{2}/(1+\varepsilon)$. Thus, by separating cases depending on whether $4\sigma_{d}^{2}\leq 1$, a short calculation gives that